OpenStudy (anonymous):

Simplify the rational expression: (m+2)/(m^2-3m-10)

OpenStudy (anonymous):

1/(m-5)

OpenStudy (anonymous):

\[\frac{m+2}{m^2-3m-10}=\frac{m+2}{(m+2)(m-5)}=\frac{1}{m-5}\]

OpenStudy (anonymous):

what eliza said, sorry

OpenStudy (anonymous):

(m-2)/(m-2)(m-5) so answer=1/(m-5)